OpenStudy (anonymous):

Complete the factoring 3y+30=3( )

OpenStudy (anonymous):

3(y+10)

OpenStudy (zepp):

If you send the 3 to the other side, it would be \[(3y + 30)/3\] 3y/3 = y 30/3 = 10 (y+10) is the answer